# Floods do not sink prices, historical memory does: How flood risk impacts the Italian housing market

### v1

@@ 4.2 Homes at risk sell at lower price in frequently flooded regions (“historical memory”)
Risk -0.029∗∗∗ -$0∗∗$008∗∗
(0.010) (0.003) (0.003)
Risk $\times$ Frequently flooded regions -0.036∗∗∗ -0.013∗∗∗ -0.014∗∗∗
(0.014) (0.004) (0.004)
Risk $\times$ Moderately flooded regions -0.042∗∗∗ -0.004 0.004
(0.013) (0.006) (0.007)
Risk $\times$ Rarely flooded regions 0.041∗∗∗ 0.003 0.001
(0.014) (0.007) (0.009)
Observations 552761 552761 552761
@@
Spatial fixed effects Municipality OMI zone Census Tract
Number of clusters 7203 20297 160630
@@
Table 2 shows the results. We only report the effect of flood risk on price, but the other hedonic coefficients for Model 3 can be found in Table LABEL:tab:hedonic. In columns 1-3-5, we show the national average effect of flood risk on home prices, whereas columns 2-4-6 disaggregate this effect across frequently, moderately, and rarely flooded regions. Models 1-2 consider spatial fixed effects at the level of the municipality, while models 3-4 and 5-6 consider OMI microzone and census tract fixed effects, respectively. For each coefficient, we report the standard error and the significance level. In all cases, standard errors are clustered at the corresponding spatial level and we use the within estimator. [^28] We tested a few specifications for the independent variables, such as whether we should consider the logarithm of the surface area, floor number as a numeric variable or binned categories, or whether we should remove outliers. In the end, we adopted the following baseline specification: logarithmic home size (m2), binned home floor levels and removing the top and bottom 0.1% samples for all numeric variables. Supplementary Section B.2 reports the robustness analysis for alternative specifications and further tests, like the robustness analyses for the definition of the historical memory. All results shown are in line with those presented in the main text.
@@
Considering the national effect (first row), Model 1, which assumes municipality-level fixed effects, shows a 2.9% price reduction for homes at risk, while Models 3 and 5, with OMI microzone and census tract fixed effects, show a 1% and 0.8% depreciation, respectively. [^29] It was unexpected that larger aggregation levels for spatial fixed effects would lead to stronger price reduction. In fact, we expected the opposite assuming that there are positive amenities associated with being located close to rivers (such as water views). These amenities would lead to a positive effect of flood risk on home prices (Bin and Kruse, 2006; Daniel et al., 2009a) when failed to be controlled for, such as when controlling for the municipality instead of more detailed OMI microzones or census tracts. Because we find a stronger reduction when controlling for the municipality, it is likely that in our sample there are on average no positive amenities for being located close to rivers. We think that this is reasonable given that we consider the entire national territory, including many rural areas. [^30] As described in Section 2.2, ISPRA divides flood-prone areas into low, medium, and high risk areas, but this classification is often inconsistent across regions. Table A5 disaggregates the effect in first row of Table 2 by risk level. In all cases, being at high risk has no statistically significant effect on prices, while being at medium or low risk has a negative and statistically significant effect on prices (but whether this effect is larger in medium or low risk areas depends on the considered spatial fixed effects). We interpret these results as evidence that the risk levels are not very informative due to idiosyncratic definitions by different river basin authorities, while an indicator of risk (irrespective of the risk level) is more likely to be informative. Further details can be found in Supplementary Section B.2. Because results are very similar for OMI microzones and census tracts, from now on we consider OMI microzones as the baseline as they are the best trade-off between granularity and sample size.
@@
Model 4 shows the disaggregation of flood risk effects across frequently, moderately, and rarely flooded regions. Notably, the statistically significant 1% price reduction at the national scale is mainly driven by the depreciation observed in frequently flooded regions, reaching 1.3%. The other coefficients for flood risk are not statistically significant and show an average 0.4% price reduction in moderately flooded regions and a 0.3% increment for homes at risk in rarely flooded regions. The leading role of frequently flooded regions in determining the impacts of flood risk on prices is robust considering more granular fixed effects (Model 6 of Table 2) and different data and definitions of regions’ historical memory (we refer the reader to Supplementary Section B.2, Tables A7 and A8, for further details).
@@
Figure 4: Effect of being at flood risk on home price, disaggregated by region. Panel a): The dots show the point estimate of the coefficients, while the error bars show 95% confidence intervals. The color code reports the region classification in frequently, moderately, and rarely flooded regions. For graphical purposes, we limited the x-axis between -0.1 and 0.1 (although not fully shown, the average price variations for Molise and Basilicata are 0.19 and 0.15, respectively). Panel b): Map of the Italian regions color-coded according to the historical memory classification. Northern regions: Piemonte, Valle d’Aosta, Liguria, Lombardia, Trentino-Alto Adige, Veneto, Friuli-Venezia Giulia, and Emilia-Romagna. Center regions: Toscana, Umbria, Marche, Lazio, Abruzzo, and Molise. Southern and Islands regions: Campania, Puglia, Basilicata, Calabria, Sicilia, and Sardegna.
@@
Further to analyzing the average effect across the entire national territory and the classification of historical memory, we also look at differential effects across regions. Focusing on the same specification as Model 3 of Table 2, and interacting the risk indicator with region dummies, Figure 4 shows that the effect of risk on prices varies a lot across regions. In Emilia-Romagna, Liguria, and Toscana, homes at risk sell at a substantially higher discount than the national average (3.5%, 2.5% and 4%, respectively, compared to 1%), whereas in Abruzzo and Calabria they sell at a premium (especially for Calabria, this may reflect an idiosyncratic definition of risk in ISPRA maps, see Section 2.2). [^31] Interestingly, the average flood risk coefficient for Emilia-Romagna agrees with $\beta_{1}$ and $\beta_{2}$ retrieved using the diff-in-diff set-up (Figure A4). These two coefficients do not need to correspond exactly to the coefficient on Emilia-Romagna in Figure 4 because they only refer to the affected municipalities and not to the entire region. However, their agreement enforces our results both for the hedonic-regression and diff-in-diff analyses. In most other regions the effects are not statistically significant at the 5% level, although regions in the North and Center are more likely to show negative effects than regions in the South and Islands. Moreover, four out of seven frequently flooded regions present price reduction due to flood risk, while rarely flooded regions exhibit price increments (5/6). We do not see a clear trend in moderately flooded regions.
@@ 4.3 Differential effects across socio-economic groups
The results presented in the previous sections support our hypothesis that flood-risk driven home price reductions are due to the historical memory of previous floods and not to specific events. Here, we investigate whether buyers with different socio-economic status (i.e., age and income) obtain different discounts when buying at-risk homes.
@@
We start by analyzing the differences in home prices considering buyers’ age (as proxied by under-36 mutuo giovani, see Section 3.4). Figure 5 shows the discount obtained by Young and NotYoung buyers when purchasing homes at risk, distinguishing between rarely, moderately and frequently flooded regions. In rarely flooded regions, Young and NotYoung buyers do not get any statistically significant price change for homes in flood-prone areas. As flood frequency increases, NotYoung buyers exhibit higher and higher price reductions, reaching 2% in frequently flooded regions. Conversely, we do not find a clear trend for Young buyers: they buy homes in risky areas at a premium in moderately flooded regions but no significant effect is observed in frequently flooded regions. We interpret this result as evidence that experienced buyers (which might have witnessed several flood events in their lifetime) pay more attention to flood risk while purchasing a home.
@@
Figure 5: Flood risk effect on home prices disaggregated across historical memory regions (rows) and mortgage applicant’s age, Equation (4). The age is proxied considering whether the transaction is financed by a mutuo giovani (red dot) or not (blue square).
@@
Building on the results presented in Figure 5, we deepen our analysis by considering also the income class of the buyers, Equation (5). Figure 6 summarizes our findings. Focusing on the Young buyers (left panel), only high income buyers in moderately and frequently flooded regions obtain a statistically significant price reduction (around 2%) when purchasing a home in areas at risk. Instead, low- and medium-income buyers purchase homes in flood-prone areas at a premium. Instead, NotYoung buyers (right panel) in all income classes obtain a significant discount (around 1.5%) for at-risk homes in frequently flooded regions. This behavior holds for the moderately flooded regions when medium and high income buyers are considered. In rarely flooded regions, instead, no buyers of any age or income class exhibit significant price changes when purchasing at-risk properties. While these results could be explained by linking income to education (Bradt et al., 2021), we leave the exploration of this hypothesis to future work since it requires additional data (such as precise the education level of the buyers) that we lack.
@@
Figure 6: Flood risk effect on home prices disaggregated across flood frequency regions, mortgage applicant age, and income class of the buyer, Equation (5). The left panel shows the home price variation due to flood risk for transactions financed by the mutuo giovani, and disaggregated by our classification of historical memory and buyers’ income class. The right panel reports the same analysis for transaction financed by a standard mortgage. The buyers’ income class is estimated considering the top, middle and bottom terciles of the income distribution within the same region (see Section 3.4 for further details).
@@
Table 3 reports the effect of flood risk on buyers’ income, estimated following Equation (6). We study the risk coefficient on incomes, even including the transaction price as an independent variable. If including the price as a dependent variable makes any effect on income non-significant, it suggests that lower prices are sufficient to explain why lower-income households sort into at-risk areas. Vice versa, other factors such as lack of information on flood risk for lower-income households may be at play.
Table 3: Flood risk effect on buyers’ incomes, Equation (6). Models 1-2 report the effect of the risk on buyers’ income at the national scale (Model 1) and disaggregated across regions with different historical memory (Model 2). Models 3-4 show the same analyses, including the logarithm of the home sale price as an independent variable. *** $p<0.01$, ** $p<0.05$, * $p<0.1$.
@@
Risk -0.009∗∗∗ -0.005
(0.003) (0.003)
Risk $\times$ Frequently flooded regions -0.013∗∗∗ -0.008∗∗
(0.004) (0.003)
Risk $\times$ Moderately flooded regions -0.004 -0.003
(0.006) (0.006)
Risk $\times$ Rarely flooded regions 0.009 0.007
(0.008) (0.008)
Observations 552761 552761
@@
Spatial fixed effects OMI zone OMI zone
Number of clusters 20297 20297
@@
Models 1-2 do not consider the price as an independent variable. On average across the whole Italian territory (Model 1), people buying a home in a flood-risk area have an income 0.9% lower than people living in non-risk areas. Model 2 highlights that buyers’ income differences are larger than the national average (and significant) for the frequently flooded regions while non-statistically significant income variation are observed in moderately and rarely flooded regions. Once the price correction is applied (Models 3-4), the Italian average risk coefficient (Model 3) increases to 0.5% and remains statistically significant at 10% level. Disaggregating the national average according to regions’ historical risk classification (Model 4), we note that people buying homes in risk-prone areas are 0.8% poorer than those purchasing homes in safe zones in frequently flooded regions. This result suggests that people with lower income mostly buy homes at risk because they are cheaper, but other factors related to the flood risk contribute to explaining the income differences.
@@ 5 Conclusion
This is the first work that systematically analyzes the impact of flood risk on home prices throughout Italy. Our study reveals three key findings. First, unlike much of the existing literature, we find that even the largest recent flood in Italy–the 2023 Emilia-Romagna flood–did not lead to additional price declines for at-risk homes that were not directly hit. Second, while homes in flood-prone areas generally sell at a small discount of about 1%, this penalty is concentrated in regions with frequent flooding, where it increases up to 4%. Third, younger buyers, who have experienced fewer floods in their lifetime, do not obtain a discount when purchasing at-risk properties, whereas more experienced buyers do. Moreover, lower-income buyers are more likely to settle in flood-prone areas, particularly in the most frequently flooded regions, suggesting a socioeconomic sorting effect.
To make sense of these findings, we hypothesize that it is the repeated exposure to floods, rather than isolated events, what ultimately leads to the capitalization of flood risk in home prices. We argue that, with no institutional requirements for flood risk disclosure or insurance that encourages the sharing of information regarding flood risk, households rely on the historical memory of major floods to identify the most at-risk areas. This explains why price penalties emerge only in regions with frequent floods and why experienced buyers–but not younger ones–obtain price discounts for at-risk properties. Finally, we find that lower-income buyers are more likely to settle in flood-prone areas, highlighting potential socioeconomic disparities in climate risk exposure.

### v2

@@ 4.2 Homes at risk sell at lower price under high awareness conditions
Risk
-0.029∗∗∗
-0.010∗∗∗
-0.008∗∗
Risk $\times$ High
awareness
-0.038∗∗
-0.017∗∗∗
-0.012∗∗∗
Risk $\times$ Medium
awareness
-0.031∗∗∗
-0.006
-0.007
Risk $\times$ Low
awareness
-0.015
-0.002
-0.005
Observations 552728 552728 552728
@@
Spatial fix. effects Municipality OMI zone Census Tract
N. clusters 7203 20297 160630
@@
Table 2 shows the results. We only report the effect of flood risk on price, but the other hedonic coefficients can be found in Table LABEL:tab:hedonic (using Model 3 as a reference). In columns 1-3-5, we show the national average effect of flood risk on home prices, whereas columns 2-4-6 disaggregate this effect according to the level of awareness at the time of sale (high, medium or low awareness). Models 1-2 consider spatial fixed effects at the level of the municipality, while models 3-4 and 5-6 consider OMI microzone and census tract fixed effects, respectively. For each coefficient, we report the standard error and the significance level. In all cases, standard errors are clustered at the corresponding spatial level and we use the within estimator. [^34] We tested a few specifications for the independent variables, such as whether we should consider the logarithm of the surface area, floor number as a numeric variable or binned categories, or whether we should remove outliers. In the end, we adopted the following baseline specification: logarithmic home size (m2), binned home floor levels and removing the top and bottom 0.1% samples for all numeric variables. Supplementary Section B.2 reports the robustness analysis for alternative specifications and further tests, like the robustness analyses across alternative half times ($\tau=7$ and $17$ years) and event records. All results shown are in line with those presented in the main text.
@@
Considering the national effect (first row), Model 1, which assumes municipality-level fixed effects, shows a 2.9% price reduction for homes at risk, while Models 3 and 5, with OMI microzone and census tract fixed effects, show a 1% and 0.8% depreciation, respectively. [^35] It was unexpected that larger aggregation levels for spatial fixed effects would lead to stronger price reduction. In fact, we expected the opposite assuming that there are positive amenities associated with being located close to rivers (such as water views). These amenities would lead to a positive effect of flood risk on home prices (Bin and Kruse, 2006; Daniel et al., 2009a) when failed to be controlled for, such as when controlling for the municipality instead of more detailed OMI microzones or census tracts. Because we find a stronger reduction when controlling for the municipality, it is likely that in our sample there are on average no positive amenities for being located close to rivers. We think that this is reasonable given that we consider the entire national territory, including many rural areas. [^36] As described in Section 2.2, ISPRA divides flood-prone areas into low, medium, and high risk areas, but this classification is often inconsistent across regions. Table A5 disaggregates the national effect of flood risk on prices, shown in the first row of Table 2, by risk level. In all cases, being at high risk has no statistically significant effect on prices, while being at medium or low risk has a negative and statistically significant effect on prices (but whether this effect is larger in medium or low risk areas depends on the considered spatial fixed effects). We interpret these results as evidence that the risk levels are not very informative due to idiosyncratic definitions by different river basin authorities, while an indicator of risk (irrespective of the risk level) is more likely to be informative. Further details can be found in Supplementary Section B.2. Because results are very similar for OMI microzones and census tracts, from now on we consider OMI microzones as the baseline as they are the best trade-off between granularity and sample size.
@@
Model 4 shows the disaggregation of flood risk effects when considering our awareness measure at the time of sale. Notably, the statistically significant 1% price reduction at the national scale is mainly driven by the discount observed in high awareness conditions (-1.7%). The other coefficients for flood risk are not statistically significant and show an average 0.6% and 0.2% price reduction for transactions with medium and low awareness. The leading role of high awareness transactions in determining the impacts of flood risk on prices is robust considering more granular fixed effects (Model 6 of Table 2), different event data and half-life parameter (we refer the reader to Supplementary Section B.2, Tables A7 and A8, for further details).
@@
Figure 4: Effect of being at flood risk on home price, disaggregated by region. Panel a): The dots show the point estimate of the coefficients, while the error bars show 95% confidence intervals. For graphical purposes, we limited the x-axis between -0.1 and 0.1 (although not fully shown, the average price variations for Molise and Basilicata are 0.19 and 0.15, respectively). Regions are sorted according to time-averaged awareness: the first region (Emilia-Romagna) shows the largest average awareness value while the last region shows the lowest value. Panel b): Map of the Italian regions colored according to the average awareness at time of sale ($\tau=10$ years, see Section 2.4). Northern regions: Piemonte, Valle d’Aosta, Liguria, Lombardia, Trentino-Alto Adige, Veneto, Friuli-Venezia Giulia, and Emilia-Romagna. Center regions: Toscana, Umbria, Marche, Lazio, Abruzzo, and Molise. Southern and Islands regions: Campania, Puglia, Basilicata, Calabria, Sicilia, and Sardegna.
@@
To provide more insights into the spatial heterogeneity of flood risk we also disaggregated the national estimates across regions. Focusing on the same specification as Model 3 of Table 2, and interacting the risk indicator with region dummies, Figure 4 shows that the effect of risk on prices varies a lot across regions. In Emilia-Romagna, Liguria, and Toscana, homes at risk sell at a substantially higher discount than the national average (3.5%, 2.5% and 4%, respectively, compared to 1%), whereas in Abruzzo and Calabria they sell at a premium (especially for Calabria, this may reflect an idiosyncratic definition of risk in ISPRA maps, see Section 2.2). [^37] Interestingly, the average flood risk coefficient for Emilia-Romagna agrees with $\beta_{1}$ and $\beta_{2}$ retrieved using the diff-in-diff set-up (Figure A4). These two coefficients do not need to correspond exactly to the coefficient on Emilia-Romagna in Figure 4 because they only refer to the affected municipalities and not to the entire region. However, their agreement reinforces our results both for the hedonic-regression and diff-in-diff analyses. In most other regions the effects are not statistically significant at the 5% level, although regions in the North and Center are more likely to show negative effects than regions in the South and Islands. When comparing these results with the average awareness at the time of sale for each region, we can observe some relevant correlations. Four out of five regions in the top 25% awareness present price reduction due to flood risk (Emilia-Romagna, Liguria, Toscana, and Piemonte), while regions in the bottom 25% exhibit price increments (Marche, Basilicata, Molise, and Puglia).
@@ 4.3 Differential effects across socio-economic groups
The results presented in the previous sections support our hypothesis that flood-risk-driven home price reductions are due to the historical memory of previous floods and not to specific events. Here, we investigate whether buyers with different socio-economic status (i.e., age and income) obtain different discounts when buying at-risk homes.
@@
We start by analyzing the differences in home prices considering buyers’ age (as proxied by under-36 mutuo giovani, see Section 3.4). Figure 5 shows the discount obtained by Young and NotYoung buyers when purchasing homes at risk, distinguishing by the level of flood awareness (high, medium and low) at the time of sale. Under low awareness conditions, Young and NotYoung buyers do not get any statistically significant discount for homes in flood-prone areas. As flood awareness gets higher, NotYoung buyers exhibit higher and higher discounts, reaching 2% for transactions of risky homes in high awareness times. Conversely, we do not find a clear trend for Young buyers, with no statistically significant discounts under any awareness level. We interpret this result as evidence that experienced buyers, who may have witnessed several flood events in their lifetime, pay more attention to flood risk when purchasing a home. This is because other differences between buyers in their ability to negotiate discounts are unlikely to vary so markedly based on flood risk awareness.
@@
Figure 5: Flood risk effect on home prices disaggregated across flood awareness (rows) and mortgage applicant’s age, Equation (5). The age is proxied considering whether the transaction is financed by a mutuo giovani (red dot) or not (blue square).
@@
Building on the results presented in Figure 5, we deepen our analysis by considering also the income class of the buyers, following Equation (6). Figure 6 shows our findings. Though on average Young buyers exhibit no significant price reduction for buying homes at risk, we observe that this behavior changes when disaggregating by income levels. Young buyers (left panel) with high-income obtain a statistically significant price reduction (around 3.5%) when purchasing at-risk homes under high and low awareness conditions. This behavior is not observed in the other income levels, as low- and medium-income buyers purchase homes in flood-prone areas at a premium, ranging from 0.1-2% depending on the awareness level. The influence of income is not so marked for the NotYoung buyers (right panel). In this case, all income classes have a similar trend that follows the income-aggregated average from Figure 5. Interestingly, all income levels achieve a statistically significant discount under high awareness conditions (around 2.5%), but this effect vanishes with low awareness.
Overall, NotYoung buyers seem to be more influenced by the level of awareness all income levels than Young buyers. This further supports our hypothesis that flood risk is engraved in the historical exposure to prior floods. Conversely, other socio-economical factors, like income, are key determinants for transactions performed by Young buyers. In fact, only Young buyers with high-income purchase homes at risk with significant price reductions, while other income classes pay a premium for risky homes. Further research is required to explore the nexus among income, education, and perception of flood risk (and more in general climate-related risks). However, we leave the exploration of this hypothesis to future work since it requires additional data (such as the precise education level of the buyers) that we lack.
@@
Figure 6: Flood risk effect on home prices disaggregated across awareness, mortgage applicant age, and income class of the buyer, Equation (6). The left panel shows the home price variation due to flood risk for transactions financed by the mutuo giovani, and disaggregated by our classification of historical memory (i.e., awareness) and buyers’ income class. The right panel reports the same analysis for transactions financed by a standard mortgage. The buyers’ income class is estimated considering the top, middle and bottom terciles of the income distribution within the same region (see Section 3.4 for further details).
@@
Table 3 reports the effect of flood risk on buyers’ income, estimated following Equation (7). We study the risk coefficient on incomes, even including the transaction price as an independent variable. If including the price as a dependent variable makes any effect on income non-significant, it suggests that lower prices are sufficient to explain why lower-income households sort into at-risk areas. Vice versa, other factors such as a lack of information on flood risk for lower-income households may be at play.
Table 3: Flood risk effect on buyers’ incomes, Equation (7). Models 1-2 report the effect of the risk on buyers’ income at the national scale (Model 1) and disaggregated by the level of flood awareness at the time of the purchase (Model 2, with half life time equal to 10 years). Models 3-4 show the same analyses, including the logarithm of the home sale price as an independent variable. *** $p<0.01$, ** $p<0.05$, * $p<0.1$.
@@
Risk
-0.009∗∗∗
-0.005
Risk $\times$ High
awareness
-0.015∗∗∗
-0.008∗∗
Risk $\times$ Medium
awareness
-0.011∗∗
-0.009∗∗
Risk $\times$ Low
awareness
0.002
0.003
Observations 552728 552728
@@
Spatial fix. effects OMI zone OMI zone
N. clusters 20297 20297
@@
Models 1-2 do not consider the price as an independent variable. On average across the whole Italian territory (Model 1), people buying a home in a flood-risk area have an income 0.9% lower than people buying properties in non-risk areas. Model 2 highlights that buyers’ income differences are larger than the national average (and significant) for high awareness transactions while this effect becomes smaller for medium and low awareness transactions. Once the price correction is applied (Models 3-4), the Italian average risk coefficient (Model 3) decreases to -0.5%. This result suggests that people with lower income mostly buy homes at risk because they are cheaper, but other factors related to the flood risk contribute to explaining the income differences.
@@ 5 Conclusion
This is the first work that systematically analyzes the impact of flood risk on home prices throughout Italy. Our study reveals three key findings. First, unlike much of the existing literature, we find that even the largest recent flood in Italy–the 2023 Emilia-Romagna flood–did not lead to additional price declines for at-risk homes that were not directly hit. Second, while homes in flood-prone areas generally sell at a small discount of about 1%, this penalty is concentrated in regions with high awareness conditions (i.e., frequently flooded), where at-risk homes sell at up to 4% less than safe homes. Third, experienced buyers obtain a discount when purchasing at-risk properties, whereas only high-income young buyers do so. Low- and middle-income young buyers purchase at-risk homes at premium. Moreover, lower-income buyers are more likely to settle in flood-prone areas, particularly in the regions with high awareness, suggesting a socioeconomic sorting effect.
To make sense of these findings, we hypothesize that it is the repeated exposure to floods, rather than isolated events, what ultimately leads to the capitalization of flood risk in home prices. We argue that, with no institutional requirements for flood risk disclosure or insurance that encourages the sharing of information regarding flood risk, households rely on the historical memory of major floods to identify the most at-risk areas. This explains why price penalties emerge only for transactions under high awareness conditions and why experienced buyers–but not younger ones–obtain price discounts for at-risk properties. Finally, we find that lower-income buyers are more likely to settle in flood-prone areas, highlighting potential socioeconomic disparities in climate risk exposure, especially for the next years.